find L^-1 {2s+4 / s(s^2+4)}

2s+4 Find L s(s2+4) 5 -30 (write 576 by 6 e^{-3t} by e and sin(2t) or cos(3t) by sin(2t) or cos(3t).
Answer #1

doubt or problem in this then comment below .. i will help you ..

.

please thumbs up for this solution .. thanks ..

.

\frac{2s+4}{s(s^{2}+4)} = \frac{1}{x} + \frac{2-s}{s^2+4}

.

so its laplace inverse is

= 1 + sin(2t) - cos(2t)